# A Comprehensive Library for Benchmarking Multi-class Visual Anomaly Detection

### v4

@@ Appendix I Detailed Quantitative Results on MPDD Dataset Under 100 epochs
Table A57: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
@@
Table A58: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
@@
Table A59: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
@@
Table A60: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
@@
cable_gland Aug. DRAEM [37] 26.1 68.5 89.2 56.9 00.4 00.9 26.6 00.5 38.4
@@
Table A61: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
@@
Table A62: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
@@
Table A63: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
@@
juice_bottle Aug. DRAEM [37] 52.9 75.6 83.4 40.0 05.3 10.2 14.7 05.4 40.3
@@
Table A64: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
@@
[…]
splicing_connectors Aug. DRAEM [37] 44.0 56.7 76.4 50.5 07.4 14.1 17.5 07.6 38.1
@@
Table A65: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
@@
Table A66: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
@@
Table A67: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
@@
pipe_fryum Aug. DRAEM [37] 63.2 76.9 80.8 46.0 01.3 03.1 14.3 01.6 40.8
@@
Table A68: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
@@
Table A69: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
@@
Table A70: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
@@
[…]
bracket_black Aug. DRAEM [37] 45.3 54.3 77.0 52.8 00.2 00.5 24.9 00.3 36.4
@@
bracket_brown Aug. DRAEM [37] 53.9 69.1 80.3 54.4 00.7 01.3 28.8 00.6 41.2
@@
bracket_white Aug. DRAEM [37] 69.4 70.9 74.0 66.1 00.1 00.2 23.6 00.1 43.5
@@
Table A71: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
@@
[…]
metal_plate Aug. DRAEM [37] 71.7 89.7 84.5 57.5 30.0 21.2 25.7 11.8 54.3
@@
Table A72: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A73: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A74: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
@@ Appendix V Detailed Quantitative Results on MAD_Sim Dataset Under 300 epochs
Table A75: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A76: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A77: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A78: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A79: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A80: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.

### v6

@@ Appendix I Detailed Quantitative Results on MPDD Dataset Under 100 epochs
Table A57: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. $3.$
@@
Table A58: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. $3.$
@@
Table A59: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. $3.$
@@
Table A60: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. $3.$
@@
cable_gland
Aug.
DRAEM [37] 26.1 68.5 89.2 56.9 00.4 00.9 26.6 00.5 38.4
@@
Table A61: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. $3.$
@@
Table A62: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. $3.$
@@
Table A63: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. $3.$
@@
juice_bottle
Aug.
DRAEM [37] 52.9 75.6 83.4 40.0 05.3 10.2 14.7 05.4 40.3
@@
Table A64: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. $3.$
@@
mAU-ROC mAP $F_{1}-m$ax mAU-ROC mAP $F_{1}-m$ax
splicing_connectors
Aug.
DRAEM [37] 44.0 56.7 76.4 50.5 07.4 14.1 17.5 07.6 38.1
@@
Table A65: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. $3.$
@@
Table A66: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. $3.$
@@
Table A67: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. $3.$
@@
pipe_fryum
Aug.
DRAEM [37] 63.2 76.9 80.8 46.0 01.3 03.1 14.3 01.6 40.8
@@
Table A68: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. $3.$
@@
Table A69: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. $3.$
@@
Table A70: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. $3.$
@@
mAU-ROC mAP $F_{1}-m$ax mAU-ROC mAP $F_{1}-m$ax
bracket_black
Aug.
DRAEM [37] 45.3 54.3 77.0 52.8 00.2 00.5 24.9 00.3 36.4
@@
bracket_brown
Aug.
DRAEM [37] 53.9 69.1 80.3 54.4 00.7 01.3 28.8 00.6 41.2
@@
bracket_white
Aug.
DRAEM [37] 69.4 70.9 74.0 66.1 00.1 00.2 23.6 00.1 43.5
@@
Table A71: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. $3.$
@@
mAU-ROC mAP $F_{1}-m$ax mAU-ROC mAP $F_{1}-m$ax
metal_plate
Aug.
DRAEM [37] 71.7 89.7 84.5 57.5 30.0 21.2 25.7 11.8 54.3
@@
Table A72: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A73: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A74: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. $3.$
@@ Appendix V Detailed Quantitative Results on MAD_Sim Dataset Under 300 epochs
Table A75: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A76: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A77: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A78: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A79: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A80: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. $3.$